# Efficient electronic cooling above 2 K by niobium-based superconducting tunnel junctions

### v1

@@ Abstract
Numerous applications, from industrial non-destructive imaging through ultra-sensitive photon counting to various implementations of solid-state quantum computers require low temperatures for their sensor and processor chips. Replacing the bulky cryo-liquid based cooling stages of cryo-enabled instruments by chip scale refrigeration is envisioned to disruptively reduce the system size similarly as microprocessors did for computers. Chip scale cooling has been demonstrated with electronic refrigerators based on tunnel junctions in the sub-1 K temperature range. Here, we extend the operation temperature to above 2 K, thereby, bridging the gap between electronic and pulse tube refrigerators. We report on scalable Al–AlOx–Nb superconducting tunnel junction cooler technology that can deliver electronic cooling power of mW/mm, which is enough to demonstrate significant electron temperature reduction of 0.82 K against the phonon bath at 2.4 K (34% relative cooling). Our work shows that the key material of integrated superconducting circuits–niobium–enables powerful cryogenic refrigerator technology. This result is a prerequisite for practical cryogenic chip scale refrigerators and, at the same time, it introduces a new electro-thermal tool for quantum heat transport experiments.
@@
Figure 1 displays device micrographs, schematics of the experimental setting and heat flow schematics. In Figs. 1a) and b), respectively, we show a false-colored scanning electron micrograph of the refrigerator with a simplified experimental electronic circuit schematics (see supplementary section S1 for more details) and false-colored cross-section transmission electron micrographs of the tunnel junction. The device has a 20 nm thick Al film on oxidized Si substrate common to four junctions connected to it. Two larger junctions (2.8 m each, referred as cooler later in the text) on the left and the right of the refrigerator are used for cooling; the smaller junction on the top (0.49 m) is used as a thermometer and the symmetric junction on the bottom provides an independent probe for the electric potential of the Al electrode. The refrigerator was fabricated on a 150 mm p-type Si wafer with 200 nm thick thermal SiO, using a variant of side-wall spacer passivated structure (SWAPS) fabrication process, which was originally developed for Nb–Nb Josephson junctions [33], and has been used, e.g., for the scalable fabrication of Josephson parametric amplifiers [34] and qubits [35]. A trilayer with Al (20 nm)–AlOx- Nb (60 nm) was deposited in-situ with direct current (DC) sputtering and patterned with ultraviolet stepper lithography and reactive ion etching. A side-wall spacer passivation layer of SiO was deposited by plasma-enhanced chemical vapor deposition (PECVD) and plasma etched to form the SWAPS [33] to prevent direct galvanic contacts from the bottom Al electrode on the trilayer to the wiring layer of Nb (100 nm) - Al (30 nm) deposited on top. SWAPS spacer and device cross-section are displayed in the panels of Fig. 1c). Prior to depositing the wiring layer, argon milling was performed to ensure a galvanic contact between the trilayer top Nb and the Nb wiring layer. The 30 nm thick Al top-layer was preventively designed to act as a quasiparticle trap to enhance the cooler performance [36, 37]. An important figure of merit in the quality of superconducting tunnel junction processes is the normalized sub-gap leakage, typically parametrized through the phenomenological Dynes parameter [38]. A set of simple cross-wire tunnel junctions was fabricated with a dedicated diagnostic stack (see supplementary section S2). Based on measurements of their differential conductance, the typical superconducting energy gap in Nb was found to be meV in the operative range of interest, with relative sub-gap leakage lower than 10, ideally suited to electron cooling applications.
@@
Figure 1c) shows the thermal model for the refrigerator. The Al electron gas is cooled by thermionic energy filtering of the tunneling electrons through AlOx to Nb [20]. The electron system is coupled to the Al phonon system, resulting in -dependent electron-phonon heat flow when Al is in the normal state [20]. The Al phonon bath is further connected to the silicon chip and the sample stage at the cold head of the cryostat, but due to the large contact areas, all aforementioned phonon baths are assumed to be in equilibrium. Thus the relevant heat flows in the refrigerator are from the electron cooling and the counteracting electron-phonon coupling. The temperature of the Al electron system was determined by junction thermometry (see below).
@@
The refrigerator was characterized in a adsorption cryostat first by DC biasing the cooler junctions and recording the corresponding voltage to find the regions of interest in the current-voltage (IV) characteristics (see supplementary section S1). The junctions were biased with current sources due to the low resistance of the junctions (from below 1 to 1.5 k as a function of temperature) compared to the line resistances in the cryostat ( 0.3 k per line). Cooling measurements were performed by DC biasing the cooler at points of interest, and recording the thermometer characteristics by varying the amplitude of a zero-to-positive square wave (frequency: 200 Hz) current bias and deriving the corresponding voltage from the first harmonic of a lock-in readout [39]. The method allows measuring sub- signals with intrinsic offset-zeroing (rejecting the slow voltage preamplifier drift), as discussed in the supplementary section S1 in more detail.
@@
IV characteristics at several bath temperatures are shown in Fig. 2a) for the cooler junctions up to = 3 mV and in Fig. 2b) for the thermometer junction near zero-bias (z.b.). The specific resistance of the junctions was measured to be around 17 m. At the bath temperature = 2.8 K, the cooler IV characteristics resembles that of a NIS junction, but at lower the IVs start to develop two features characterized by negative differential resistance (NDR) between cooler voltages = 2.5 mV and 3 mV. We note that the cooler is a series of two Nb-AlOx-Al junctions, where the thin Al is superconducting below 1.38 K (see supplementary section S3).
@@
Figure 4 captures the main result of the paper. There the extracted values of the thermometer at best found cooler bias values are shown as a function of on the left y-axis. The thermometer calibration was used to convert the values to the temperature which is shown on the right-side y-axis of Fig. 4a). We see cooling to below 1.6 K from all investigated values. Maximum absolute cooling is observed from = 2.4 K to 1.57 K corresponding to a relative cooling of 34% from the phonon bath (see 4b)). Importantly, through optimal cooler junction bias we can also observe the appearance of a sizeable ( 100 nA) superconducting branch at bath temperature values as high as 1.6 K, as well as enhance by more than 400% the supercurrent branch present at = 1.4 K. These effects, shown in the supplementary Fig. S7, are consistent with and corroborate the above interpretation of cooling based on the z.b. resistance thermometry.
@@
The refrigerator was simulated with a finite element method to estimate the temperature at the cooler junctions and the corresponding cooling power density (see supplementary section S6). We assumed normal state Al in calculating the thermal conductivity and electron-phonon coupling. The simulations show that the temperature across the 20 nm thick Al film is uniform up to 1.5%. The estimated cooling power density is in the order of mW/mm. In comparison, in Ref. [19] the authors report W/mm in a V-based refrigerator operating at 1 K and in Ref. [37] the reported estimate for an Al-based refrigerator operating at 0.3 K is W/mm. The cooling power of superconducting tunnel junction refrigerator scales both as a function of temperature and the energy gap. In the optimal case, the gap is matched for the operating temperature (), which results in cooling power density [25, 27]. An approximative comparison for the technical performance of different of cooler junctions operating at different temperatures can thus be obtained by comparing the values of for each case, which yields about 8 W/mmK, 65 W/mmK, and 280 W/mmK for the above-mentioned Al- and V-based devices, and our refrigerator operating at 2.4 K, respectively. This indicates that our cooler is efficient compared to other electronic refrigerators even if the larger temperature and superconducting gap are taken into account. This is also an indication of high-quality junctions in terms of transparency and leakage current, achieved with our foundry compatible Nb-based fabrication process.
@@
In conclusion, we have demonstrated significant cooling of free electrons with respect to lattice phonons up to 2.4 K bath temperature. At 2.4 K, we observe cooling of electrons down to 1.6 K – equivalent to 34% relative cooling. The temperature range was limited by the electron thermometry, i.e., the cooling effect is expected to work similarly also above 2.4 K. The cooling result was enabled by developing Nb-based superconducting tunnel junction refrigerator technology. With our modern foundry compatible Al–AlOx–Nb tunnel junctions, we obtained a cooling power density of mW/mm, which is expected to enable substantial cooling in practical chip scale refrigerators. Most importantly, after decades of research on tunnel junction based electronic coolers, here we demonstrated a device that can operate at the base temperature of pulse tube refrigerators, thus paving the way for replacing bulky dilution refrigerators by miniaturized electronic coolers that operate without rare and expensive .
@@ Acknowledgements
The authors have no conflicts to disclose.
@@
Research was conceptualized by J.H, A.R, A.K, J.L and M.P. Device design was done by J.H, A.R, R.P.L and K.V. Device fabrication was led by R.P.L with the support from J.H. T.R did device validation through SEM imaging. Device characterization was done in room temperature by J.H and J.Ge, and in low temperature by J.H and A.R. Data was analyzed by J.H, A.R, E.M and A.K. Finite element method simulation was done by E.M and A.R. with the support from J.H. Al–AlOx–Nb junction development was done by R.P.L, M.R, J-P.K, O.P, V.V, W.F, E.T.M, M.K, M.P and J.G, who also supervised the junction development. M.P supervised the refrigerator work. J.H and M.P wrote the manuscript with contributions from all other authors.

### v2

# Efficient electronic cooling by niobium-based superconducting tunnel junctions

@@ Abstract
Replacing the bulky cryo-liquid based cooling stages of cryo-enabled instruments by chip scale refrigeration is envisioned to disruptively reduce the system size similarly as microprocessors did for computers. Electronic refrigerators based on superconducting tunnel junctions have been envisioned to provide a solution, but reaching the necessary above 1 kelvin operation temperature range has remained a goal out of reach for several decades. We show efficient electronic refrigeration by Al–AlOx–Nb superconducting tunnel junctions starting from bath temperatures above 2 K. The junctions can deliver electronic cooling power up to mW/mm2, which enables us to demonstrate tunnel current driven electron temperature reduction from 2.4 K to below 1.6 K (34% relative cooling) against the phonon bath. Our work shows that the key material of integrated superconducting circuits–niobium–enables powerful cryogenic refrigerator technology. This result is a prerequisite for practical cryogenic chip scale refrigerators and, at the same time, it introduces a new electro-thermal tool for quantum heat transport experiments.
@@
On-chip refrigerator device
Figure 1 displays device micrographs, schematics of the experimental setting and heat flow schematics. In Figs. 1a) and c), respectively, we show a false-colored scanning electron micrograph of the refrigerator with a simplified experimental electronic circuit schematics (see supplementary section S1 for more details) and false-colored cross-section transmission electron micrographs of the tunnel junction. The device has a 20 nm thick Al film on oxidized Si substrate common to four junctions connected to it. Two larger junctions (2.8 m2 each, referred as cooler later in the text) on the left and the right of the refrigerator are used for cooling; the smaller junction on the top (0.49 m2) is used as a thermometer and the symmetric junction on the bottom provides an independent probe for the electrical potential of the Al electrode.
@@
The refrigerator was fabricated on a 150 mm p-type Si wafer with 200 nm thick thermal SiO2, using a variant of side-wall spacer passivated structure (SWAPS) fabrication process, which was originally developed for Nb–Nb Josephson junctions [33], and has been used, e.g., for the scalable fabrication of Josephson parametric amplifiers [34] and qubits [35]. A trilayer with Al (20 nm)–AlOx- Nb (60 nm) was deposited in-situ with direct current (DC) sputtering and patterned with ultraviolet stepper lithography and reactive ion etching. A side-wall spacer passivation layer of SiO2 was deposited by plasma-enhanced chemical vapor deposition (PECVD) and plasma etched to form the SWAPS [33] to prevent direct galvanic contacts from the bottom Al electrode on the trilayer to the wiring layer of Nb (100 nm) - Al (30 nm) deposited on top. SWAPS spacer and device cross-section are displayed in the panels of Fig. 1c). Prior to depositing the wiring layer, argon milling was performed to ensure a galvanic contact between the trilayer top Nb and the Nb wiring layer. The 30 nm thick Al top-layer was preventively designed to act as a quasiparticle trap to enhance the cooler performance [36, 37].
An important figure of merit in the quality of superconducting tunnel junction processes is the normalized sub-gap leakage, typically parametrized through the phenomenological Dynes parameter [38]. A set of simple cross-wire tunnel junctions was fabricated with a dedicated diagnostic stack (see supplementary section S2). Based on measurements of their differential conductance, the typical superconducting energy gap in Nb was found to be meV in the operative range of interest, with relative sub-gap leakage lower than 10-3, ideally suited to electron cooling applications.
@@
Figure 1c) shows the thermal model for the refrigerator. The electron gas in the Al electrode is cooled by thermionic energy filtering of the tunneling electrons through AlOx to Nb [20]. The electron system is coupled to the surrounding phonon bath, resulting in -dependent electron-phonon heat flow when Al is in the normal state [20]. The Al phonon bath is further connected to the silicon chip and the sample stage at the cold head of the cryostat, but due to the large contact areas, all aforementioned phonon baths are assumed to be in equilibrium. Thus the relevant heat flows in the refrigerator are from the electron cooling and the counteracting electron-phonon coupling. The temperature of the Al electron system was determined by junction thermometry (see below).
Electro-thermal performance
@@
The refrigerator was characterized in a cryostat first by DC biasing the cooler junctions and recording the corresponding voltage to find the regions of interest in the current-voltage (IV) characteristics (see supplementary section S1). The junctions were biased with current sources due to the low resistance of the junctions (from below 1 to 1.5 k as a function of temperature) compared to the line resistances in the cryostat ( 0.3 k per line). Cooling measurements were performed by DC biasing the cooler at points of interest, and recording the thermometer characteristics by varying the amplitude of a zero-to-positive square wave (frequency: 200 Hz) current bias and deriving the corresponding voltage from the first harmonic of a lock-in readout [39]. The method allows measuring sub- signals with intrinsic offset-zeroing (rejecting the slow voltage preamplifier drift), as discussed in the supplementary section S1 in more detail.
@@
IV characteristics at several bath temperatures are shown in Fig. 2a) for the cooler junctions up to = 3 mV and in Fig. 2b) for the thermometer junction near zero-bias (z.b.). The specific resistance of the junctions was measured to be around 17 m2. At the bath temperature = 2.8 K, the cooler IV characteristics resembles that of a NIS junction, but at lower the IVs start to develop two features characterized by negative differential resistance (NDR) between cooler voltages = 2.5 mV and 3 mV. We note that the cooler is a series of two Nb-AlOx-Al junctions, where the thin Al is superconducting below 1.38 K (see supplementary section S3).
@@
Figure 4 captures the main result of the paper. The extracted values of the thermometer at best found cooler bias values are shown as a function of on the right-side y-axis of Fig. 4a). The thermometer calibration was used to convert the values to the temperature which is shown on the left-side y-axis of Fig. 4a). We see cooling to below 1.6 K from all investigated values. Maximum absolute cooling is observed from = 2.4 K to 1.57 K corresponding to a relative cooling of 34% from the phonon bath (see 4b)). Importantly, through optimal cooler junction bias we can also observe the appearance of a sizeable ( 100 nA) superconducting branch at bath temperature values as high as 1.6 K, as well as enhance by more than 400% the supercurrent branch present at = 1.4 K. These effects, shown in the supplementary Fig. S7, are consistent with and corroborate the above interpretation of cooling based on the z.b. resistance thermometry.
The refrigerator was simulated with a finite element method to estimate the temperature at the cooler junctions and the corresponding cooling power density (see supplementary section S6). We assumed normal state Al in calculating the thermal conductivity and electron-phonon coupling. The simulations show that the temperature across the 20 nm thick Al film is uniform up to 1.5%. The estimated cooling power density is in the order of mW/mm2. In comparison, in Ref. [19] the authors report W/mm2 in a V-based refrigerator operating at 1 K and in Ref. [37] the reported estimate for an Al-based refrigerator operating at 0.3 K is W/mm2. Remarkably we reach cooling power similar to that of thermionic refrigerator based on asymmetric double barrier semiconductor heterostructures operating at room temperature [43].
@@
Cooling power of superconducting tunnel junction refrigerator scales both as a function of temperature and the superconductor energy gap. In the optimal case, the gap is matched for the operating temperature (), which results in cooling power density [25, 27]. An approximative comparison for the technical performance of different of cooler junctions operating at different temperatures can thus be obtained by comparing the values of for each case, which yields about 8 W/mm2K2, 65 W/mm2K2, and 280 W/mm2K2 for the above-mentioned Al- and V-based devices, and our refrigerator operating at 2.4 K, respectively. This indicates that our cooler is efficient compared to other electronic refrigerators even if the larger temperature and superconducting gap are taken into account. This is also an indication of high-quality junctions in terms of transparency and leakage current, achieved with our foundry compatible Nb-based fabrication process.
Conclusions
@@
In conclusion, we have demonstrated significant cooling of free electrons with respect to lattice phonons up to 2.4 K bath temperature. At 2.4 K, we observe cooling of electrons down to 1.6 K – equivalent to 34% relative cooling. The temperature range was limited by the electron thermometry, i.e., the cooling effect is expected to work similarly also above 2.4 K. The cooling result was enabled by developing Nb-based superconducting tunnel junction refrigerator technology. With our modern foundry compatible Al–AlOx–Nb tunnel junctions, we obtained a cooling power density of mW/mm2, which is expected to enable substantial cooling in practical chip scale refrigerators. Most importantly, after decades of research on tunnel junction based electronic coolers, here we demonstrated a device that can operate at the base temperature of pulse tube refrigerators, thus paving the way for replacing bulky dilution refrigerators by miniaturized electronic coolers that operate without rare and expensive .
@@ Acknowledgements
The authors have no conflicts of interest to disclose.
@@
Research was conceptualized by J.H, A.R, A.K, J.S.L and M.P. Device design was done by J.H, A.R, R.P.L and K.V. Device fabrication was led by R.P.L with the support from J.H. T.R did device validation through SEM imaging. Device characterization was done in room temperature by J.H and J.Ge, and in low temperature by J.H and A.R. Data was analyzed by J.H, A.R, E.M and A.K. Finite element method simulation was done by E.M and A.R. with the support from J.H. Al–AlOx–Nb junction development was done by R.P.L, M.R, J-P.K, O.P, V.V, C.F, E.T.M, M.K, M.P and J.G, who also supervised the junction development. M.P supervised the refrigerator work. J.H and M.P wrote the manuscript with contributions from all other authors.